### v4

@@ 3.3.2 Simulations
Among the methods that can be applied in this context of FOSR with functional coefficients expanded via basis functions, we can highlight LASSO (Least Absolute Shrinkage and Selection Operator) ([36]), SCAD (Smoothly Clipped Absolute Deviation) penalized regression ([9]) and MCP (Minimax Concave Penalty) penalized regression ([40]), which are regularization and variable selection methods for regression analysis. However, when any of these methods are applied to functional regression after the basis expansion of the functional coefficients, the variable selection is not guaranteed since the method will not consider the existence of groups among the scalar coefficients derived from the basis expansions. In other words, selection of some of the expansion bases may occur; however, not of the variables, since for that, it would be necessary to zero the functional coefficients as a whole, which would mean to zero together all the scalar coefficients of the expansion of the respective functional coefficient.
Naturally, when applying a procedure that takes into account the structure in groups in order to simultaneously include or exclude the groups of these scalar coefficients, one truly has a variable selection mechanism for the aforementioned functional structure. Fortunately, one can find papers in the literature of linear models that extend the ideas of [36], [9] and [40] for the context of selecting groups of variables. First, [39] proposed group LASSO (grLASSO), which performs this type of group selection, and then [37] proposed a group selection method which uses the SCAD penalty (grSCAD), while [17] showed something similar for the MCP penalty (grMCP).
Thus, the regularization and selection methods in groups described above (grLASSO, grSCAD and grMCP) are considered in this section to carry out a comparative analysis with our proposed method in regards to the fit quality and the correct selection capacity of the variables in each one of the methods in question. We used synthetic data replications generated as described in Section 3.1 with $m=10$ curves and $K=10$ B-spline bases for each functional coefficient in each replication. Finally, it is worth noting that the models are compared in two scenarios, one with low dispersion data ($\sigma=0.2$) and another with higher dispersion data ($\sigma=20$), and 100 replications were used in each of these scenarios.
The grpreg function from the grpreg package ([5]) of the statistical software R was used to obtain the results of the grLASSO, grSCAD and grMCP models. For the implementation of the estimation algorithm of the aforementioned methods, the grpreg function refers to the works presented in [4] and [17].
[…]
$(\boldsymbol{X_{p\times m}})^{{}^{\prime}}\otimes\boldsymbol{B_{n\times K}}=\left[\begin{matrix}x_{11}\boldsymbol{B}&\dots&x_{p1}\boldsymbol{B}\\ \vdots&\dots&\vdots\\ x_{1m}\boldsymbol{B}&\dots&x_{pm}\boldsymbol{B}\end{matrix}\right]\text{.}$
For the group argument of the grpreg function, one must define a vector that describes the grouping of the coefficients, whereas, for the penalty argument, one can select either “grLASSO,” “grSCAD” or “grMCP” expression to define the model to be used.
Finally, the grpreg function requires the definition of a search grid for the regularization parameter $\lambda$ or the definition of a single value for it. Thus, in this work, we chose to define a single $\lambda$ value within the function at a time. However, we ran the function for different values of $\lambda$ defined over a grid characterized by a sequence that was generated in R as follows: seq(0.001,2,length = 9) for $\sigma=0.2$, and seq(0.001,4,length = 9) for $\sigma=20$. Both sequences have the same length, but different upper bounds (2 and 4), from which points we can observe that the regularization in data with greater dispersion becomes too intense. Thus, for each data dispersion level, each $\lambda$ value of the respective sequence, and each of the hundred replications, the grpreg function was applied, and the performance metrics, (15) and (16), were subsequently calculated.
@@
Figure 7: Boxplots of the metric (15) for each competing model (grLASSO, grSCAD and grMCP), according to the data dispersion degree and the respective regularization $\lambda$ values tested.
(a) Data with $\sigma=0.2$.
(b) Data with $\sigma=20$.
Figure 8: Boxplots of the MSEs for each adversary model (grLASSO, grSCAD and grMCP), according to the data dispersion degree and the respective $\lambda$ values tested.
Figure 7 and 8 present the boxplots of the metric (15) and MSE, respectively, for each competing model according to the data dispersion degree and the respective $\lambda$ values tested. Regardless of which metric is evaluated (metric (15) or MSE), we can observe that the quality of the fit starts to decrease from a certain $\lambda$ value, indicating that the regularization becomes more and more intense and farther from the ideal if large values of $\lambda$ are used. This discrepancy is even more evident in the scenario with low data dispersion since the boxplots for low values of the regularization parameter are at a level very close to the perfect value (1 for the metric (15) and 0 for the MSE), so close that it is even difficult to visualize them in the figures, while the boxplots associated with the high values of $\lambda$ are at much worse levels of the performance metrics and with larger interquartile ranges. Figure 9 and Supplementary Figure 6 present a comparison between our proposed model and the competing grLASSO, grMCP and grSCAD methods regarding the metric (15) and MSE, respectively.
Model $\hat{\beta}_{l}(.)$ $\lambda$
0.001 0.250875 0.50075 0.750625 1.0005 1.250375 1.50025 1.750125 2
grLASSO $\hat{\beta}_{1}(.)$ 1 0.02 0.01 0.01 0.01 0.01 0.01 0.01 0.01
$\hat{\beta}_{2}(.)$ 1 0.02 0.02 0.02 0.02 0.02 0.02 0.02 0.02
$\hat{\beta}_{3}(.)$ 1 1 1 1 1 1 1 1 1
$\hat{\beta}_{4}(.)$ 1 0.05 0.05 0.05 0.05 0.03 0.03 0.02 0.02
$\hat{\beta}_{5}(.)$ 1 1 1 1 1 1 0.99 0.97 0.96
$\hat{\beta}_{6}(.)$ 1 0.09 0.07 0.06 0.04 0.04 0.04 0.04 0.03
grMCP $\hat{\beta}_{1}(.)$ 1 0 0 0 0 0 0 0 0
$\hat{\beta}_{2}(.)$ 1 0 0 0.01 0.01 0 0 0 0
$\hat{\beta}_{3}(.)$ 1 1 1 1 1 1 1 1 1
$\hat{\beta}_{4}(.)$ 1 0 0 0.01 0.01 0 0 0 0
$\hat{\beta}_{5}(.)$ 1 1 1 0.99 0.99 1 0.99 0.99 0.97
$\hat{\beta}_{6}(.)$ 1 0 0 0 0 0 0 0 0
grSCAD $\hat{\beta}_{1}(.)$ 1 0 0 0 0 0 0 0 0
$\hat{\beta}_{2}(.)$ 1 0 0.01 0 0 0 0 0 0
$\hat{\beta}_{3}(.)$ 1 1 1 1 1 1 1 1 1
$\hat{\beta}_{4}(.)$ 1 0 0.01 0 0 0 0 0 0
$\hat{\beta}_{5}(.)$ 1 1 1 1 1 1 0.99 0.99 0.97
$\hat{\beta}_{6}(.)$ 1 0 0 0 0 0 0 0 0
Table 3: Proportion of non-zero estimates for each functional coefficient among the 100 replications, according to each model configuration tested. Data with $\sigma=0.2$.
Model $\hat{\beta}_{l}(.)$ $\lambda$
0.001 0.500875 1.00075 1.500625 2.0005 2.500375 3.00025 3.500125 4
grLASSO $\hat{\beta}_{1}(.)$ 1 0.94 0.62 0.2 0.08 0.04 0.03 0.03 0.01
$\hat{\beta}_{2}(.)$ 1 0.97 0.67 0.17 0.05 0.04 0.01 0 0
$\hat{\beta}_{3}(.)$ 1 1 1 1 1 1 1 1 1
$\hat{\beta}_{4}(.)$ 1 0.97 0.76 0.27 0.09 0.04 0.04 0.03 0.02
$\hat{\beta}_{5}(.)$ 1 1 1 0.98 0.96 0.9 0.88 0.75 0.63
$\hat{\beta}_{6}(.)$ 1 0.99 0.6 0.2 0.1 0.07 0.05 0.03 0.03
grMCP $\hat{\beta}_{1}(.)$ 1 0.95 0.43 0.1 0.01 0 0 0 0
$\hat{\beta}_{2}(.)$ 1 0.96 0.47 0.04 0 0 0 0 0
$\hat{\beta}_{3}(.)$ 1 1 1 1 1 1 1 1 1
$\hat{\beta}_{4}(.)$ 1 0.94 0.53 0.08 0.01 0 0 0 0
$\hat{\beta}_{5}(.)$ 1 1 0.98 1 0.98 0.92 0.84 0.74 0.54
$\hat{\beta}_{6}(.)$ 1 1 0.49 0.04 0 0 0 0 0
grSCAD $\hat{\beta}_{1}(.)$ 1 0.95 0.48 0.11 0.03 0.02 0 0 0
$\hat{\beta}_{2}(.)$ 1 0.92 0.55 0.07 0.01 0.01 0.01 0 0
$\hat{\beta}_{3}(.)$ 1 1 1 1 1 1 1 1 1
$\hat{\beta}_{4}(.)$ 1 0.94 0.56 0.12 0 0 0.01 0.01 0.01
$\hat{\beta}_{5}(.)$ 1 1 1 1 0.97 0.91 0.85 0.73 0.59
$\hat{\beta}_{6}(.)$ 1 0.96 0.55 0.09 0.03 0.02 0.01 0.01 0.01
Table 4: Proportion of non-zero estimates for each functional coefficient among the 100 replications, according to each model configuration tested. Data with $\sigma=20$.
@@
In order to demonstrate the capacity of each competing model to select the correct covariates across different values of $\lambda$ and dispersion levels, Tables 3 and 4 present the proportion of times that each functional coefficient estimate was different from zero among the total of 100 replications. We notice that no model could perform the variable selection for the lowest regularization level ($\lambda=0.001$) for all three competing methods (all proportions were equal to one). For the other $\lambda$ values tested, the models started to select the variables. Such selection for data with low dispersion is predominantly correct from $\lambda=0.250875$, while for data with large dispersion from $\lambda=2.005$. However, the non-zero proportions in Table 4 associated with the fifth functional coefficient become lower for larger values of $\lambda$.
@@
It is noted (Table 4 and Figure 8b) that the competing models are not able to balance very well the accuracy in the selection in the scenario with greater data dispersion while preserving the quality of the fit according to the metrics under study, meaning that they need to lose fit quality and increase the regularization to gain better accuracy in the selection.
For last finding that can be obtained when evaluating Tables 3 and 4 is that the grSCAD and grMCP models stand out from the grLASSO model in the scenario with low data dispersion ($\sigma=0.2$), better selecting the variables for the different $\lambda$ values tested.
@@
Tables 5 and 6 allow a brief comparison regarding the capacity to select the correct covariates among the best configurations tested for the competing models, as well as for our proposed model. The best model configuration was chosen according to the performance metric (15). For this, the average metric value over the one hundred replications was calculated for each model configuration, and we then selected the model with the configuration that returned the highest of these averages.
$\hat{Z}_{l}\hat{\beta}_{l}(.)$ (proposed model) or $\hat{\beta}_{l}(.)$ (adversary models) Proposed model ($\mu=0.3$ and $\lambda=\sqrt{2}$) Proposed model ($\mu$ as a parameter and $\lambda=\sqrt{2}$) grLASSO ($\lambda=0.001$) grMCP ($\lambda=0.250875$) grSCAD ($\lambda=0.250875$)
$1^{\text{st}}$ functional coefficient 0 0 1 0 0
$2^{\text{nd}}$ functional coefficient 0 0 1 0 0
$3^{\text{rd}}$ functional coefficient 1 1 1 1 1
$4^{\text{th}}$ functional coefficient 0 0 1 0 0
$5^{\text{th}}$ functional coefficient 1 1 1 1 1
$6^{\text{th}}$ functional coefficient 0 0 1 0 0
@@
Table 5: Proportion of non-zero estimates for each functional coefficient among the 100 replications, according to the best tested model configuration. Data with $\sigma=0.2$.
@@
$\hat{Z}_{l}\hat{\beta}_{l}(.)$ (proposed model) or $\hat{\beta}_{l}(.)$ (adversary models) Proposed model ($\mu=0.5$ and $\lambda=\sqrt{2}$) Proposed model ($\mu$ as a parameter and $\lambda=\sqrt{2}$) grLASSO ($\lambda=0.001$) grMCP ($\lambda=0.500875$) grSCAD ($\lambda=0.001$)
$1^{\text{st}}$ functional coefficient 0.04 0.01 1 0.95 1
$2^{\text{nd}}$ functional coefficient 0.02 0.01 1 0.96 1
$3^{\text{rd}}$ functional coefficient 1 1 1 1 1
$4^{\text{th}}$ functional coefficient 0.02 0 1 0.94 1
$5^{\text{th}}$ functional coefficient 1 0.99 1 1 1
$6^{\text{th}}$ functional coefficient 0.03 0.03 1 1 1
@@
Table 6: Proportion of non-zero estimates for each functional coefficient among the 100 replications, according to the best tested model configuration. Data with $\sigma=20$.
For comparisons between models in this section, specifically in relation to our proposed model, the results obtained for the simulation study presented in Subsection 3.3.2 were used. As seen in Subsection 3.3.2, the proposed model has a low sensitivity to the hyperparameter $\mu$, which is desired for conventional hyperparameters. On the other hand, it is natural that the model has some sensitivity to the hyperparameter $\lambda$ since this is the only hyperparameter that truly plays the role of regularizer ($\lambda$ is the a priori hyperparameter of the $\tau_{ki}^{2}$’s, presented in (3) and (4)). For the proposed model to not be at a disadvantage in this comparison, the ideal would be to test more $\lambda$ values, as was done for the opponents. However, only the $\lambda=\sqrt{2}$ value was tested due to the high computational cost demanded by applying the model on one hundred replications and several model configurations for $\lambda$. Even so, the proposed model with $\lambda=\sqrt{2}$ presented great performance compared to the others, with excellent fit capacity and standing out even more in the accuracy regarding the selection of covariates.
@@
The results presented in Table 5 show that most of the compared models correctly performed the covariate selection on low dispersion data ($\sigma=0.2$), except for grLASSO, which was not able to exclude any covariate in any of the hundred replications. When evaluating the selection performance for data with greater dispersion ($\sigma=20$), we can observe in Table 6 that only the two versions of the proposed model ($\mu$ as a hyperparameter and as a parameter) were able to identify the true two variables that were used in the construction of synthetic data ($3^{\text{rd}}$ and $5^{\text{th}}$) in almost all replications.
@@
It is worth noticing that the $\lambda=\sqrt{2}$ used in the versions of the proposed model was the lowest possible value for it since uniqueness problems in the stationary distribution begin to appear for smaller values of this parameter as a result of the combination of two factors: low regularization and multicollinearity of the covariate matrix. Therefore, this was a configuration with the lowest possible regularization, and the proposed model versions under this scenario were able to provide a good fit to the data while being able to select the variables correctly. In contrast, as noted earlier, the grLASSO, grSCAD and grMCP models face difficulties reconciling accuracy in the selection and quality of fit.
(a) Data with $\sigma=0.2$.
(b) Data with $\sigma=20$.
Figure 9: Boxplots of the metric (15), according to the model used and the data dispersion degree. The “$1^{\text{st}}$ Proposed Model” is considered to be the version with $\lambda=\sqrt{2}$ and $\mu$ as a parameter, while the “$2^{\text{nd}}$ Proposed Model” is the version with $\lambda=\sqrt{2}$ and $\mu$ as a hyperparameter ($\mu=0.3$ for $\sigma=0.2$ and $\mu=0.5$ for $\sigma=20$).
Figure 9 and Supplementary Figure 6 present the results regarding the performance metric in (15) and MSE, respectively, for the best model configurations. The metric (15) was used as a reference for choosing the best model configurations to be compared in Figure 9, as it was done for Tables 5 and 6. At the same time, for Supplementary Figure 6, the MSE itself was the reference metric for defining the best model settings. Thus, the comparison via MSE becomes fair. Turning the analysis to the results from data with low dispersion ($\sigma=0.2$), regardless of the metric being analyzed, slightly better performance of the competing methods can be observed in comparison to the two versions of the proposed model, especially when using the metric (15) as a reference. As for the data with greater dispersion, the models present similar goodness of fit and estimation accuracy.
Finally, Figure 10 shows the estimates obtained in each replication for the third and fifth functional coefficients for each model and each level of data dispersion. In this case, the performance metric in (15) was used as a reference for choosing the model configurations to be compared. When $\sigma=20$, it is noticeable that there is a smaller dispersion of the estimated functional coefficients in the versions of the proposed model compared to the competing models. This observation is valid both for the third and fifth functional coefficients.
(a) $\hat{\beta}_{3}(.)$ (proposed model, $\mu=0.3$, $\sigma=0.2$).
(b) $\hat{\beta}_{3}(.)$ (proposed model, $\mu=0.5$, $\sigma=20$).
(c) $\hat{\beta}_{5}(.)$ (proposed model, $\mu=0.3$, $\sigma=0.2$).
(d) $\hat{\beta}_{5}(.)$ (proposed model, $\mu=0.5$, $\sigma=20$).
(e) $\hat{\beta}_{3}(.)$ (proposed model, $\mu$ as a parameter, $\sigma=0.2$).
(f) $\hat{\beta}_{3}(.)$ (proposed model, $\mu$ as a parameter, $\sigma=20$).
(g) $\hat{\beta}_{5}(.)$ (proposed model, $\mu$ as a parameter, $\sigma=0.2$).
(h) $\hat{\beta}_{5}(.)$ (proposed model, $\mu$ as a parameter, $\sigma=20$).
(i) $\hat{\beta}_{3}(.)$ (grLASSO, $\lambda=0.001$, $\sigma=0.2$).
(j) $\hat{\beta}_{3}(.)$ (grLASSO, $\lambda=0.001$, $\sigma=20$).
(k) $\hat{\beta}_{5}(.)$ (grLASSO, $\lambda=0.001$, $\sigma=0.2$).
(l) $\hat{\beta}_{5}(.)$ (grLASSO, $\lambda=0.001$, $\sigma=20$).
(m) $\hat{\beta}_{3}(.)$ (grMCP, $\lambda=0.250875$, $\sigma=0.2$).
(n) $\hat{\beta}_{3}(.)$ (grMCP, $\quad\lambda=0.500875$, $\sigma=20$).
(o) $\hat{\beta}_{5}(.)$ (grMCP, $\quad\lambda=0.250875$, $\sigma=0.2$).
(p) $\hat{\beta}_{5}(.)$ (grMCP, $\quad\lambda=0.500875$, $\sigma=20$).
(q) $\hat{\beta}_{3}(.)$ (grSCAD, $\lambda=0.250875$, $\sigma=0.2$).
(r) $\hat{\beta}_{3}(.)$ (grSCAD, $\lambda=0.001$, $\sigma=20$).
(s) $\hat{\beta}_{5}(.)$ (grSCAD, $\lambda=0.250875$, $\sigma=0.2$).
(t) $\hat{\beta}_{5}(.)$ (grSCAD, $\lambda=0.001$, $\sigma=20$).
Figure 10: Partial functional coefficients ($3^{\text{rd}}$ and $5^{\text{th}}$, true in red and estimated for the replications in which the coefficients were selected in black), according to the data dispersion degree and model used.
3.5 BGLSS versus proposed model
To complete the comparisons with other existing methods, here is a brief comparative analysis between the proposed model and the Bayesian Group LASSO with Spike and Slab prior (BGLSS) method. The BGLSS method ([23]) deals with the Bayesian regularization and estimation problem, taking into account the existence of groups of variables. In this case, the variables of the same group must be selected or excluded together. The BGLSS function from the R package MBSGS was used to implement BGLSS. In this case, the default function was used (also with 10000 MCMC iterations, with 5000 for burn-in). Finally, the groups of scalar coefficients ($b_{.l}$’s) were defined in the same way as was done in the previous section for the other methods (grLASSO, grSCAD and grMCP), meaning the respective scalar coefficients from this expansion end up forming a group for each expansion based on a given $\beta_{l}(.)$.
As the BGLSS function returns the entire chain, estimates can be calculated separately using the map_estimate function. Thus, we have a summary measure in accordance with the one used in the other methods (Maximum a Posteriori - MAP). The chains obtained in certain replications for certain coefficients can be sparse, and this can make using the map_estimate function unfeasible, but this is not a problem in this case since there are many values numerically equal to zero and, therefore, the MAP estimate is zero itself. A relevant fact to be highlighted is that the simulated data used in this study are the same used in the previous comparative study.
(a) Data with $\sigma=0.2$.
(b) Data with $\sigma=20$.
Figure 11: Boxplots of the metric (15), according to the model used and the data dispersion degree. The “$1^{\text{st}}$ Proposed Model” is considered to be the version with $\lambda=\sqrt{2}$ and $\mu$ as a parameter, while the “$2^{\text{nd}}$ Proposed Model” is the version with $\lambda=\sqrt{2}$ and $\mu$ as a hyperparameter ($\mu=0.3$ for $\sigma=0.2$ and $\mu=0.5$ for $\sigma=20$).
Figure 11 and Supplementary Figure 7 show the boxplots for each of the performance metrics (metric (15) and MSE, respectively). In one replication with low dispersion, the BGLSS method had its worst performance resulting in an outlier in the BGLSS boxplots (for both evaluated metrics and only for $\sigma=0.2$). Figure 11 and Supplementary Figure 7 omit this outlier since we adjusted the scale to focus on the interquartile range region to enable the comparison between the boxplots. In terms of the goodness-of-fit metric (15), as shown in Figure 11, the BGLSS method can remain competitive. It is worth highlighting that the methods present a reasonably equivalent performance in this regard and that although the competing methods perform better than the two versions of the model proposed in the scenario with the low dispersion data ($\sigma=0.2$), this difference is negligible ($<0.02\%$).
@@ 3.5 BGLSS versus proposed model
Table 7 demonstrates the variable selection capacity of the BGLSS method, presenting the proportions calculated as the number of times that each functional coefficient estimate was different from zero among the 100 replications. Unlike the grLASSO, grSCAD and grMCP methods, the BGLSS method can select the variables well even in the study with more data dispersion. It is the only one among the studied models that can compete with our proposed methodology in terms of predictive quality (through the metric (15) and MSE) and variable selection capacity.
$\hat{\beta}_{l}(.)$ $\sigma$
$0.2$ $20$
$\hat{\beta}_{1}(.)$ 0 0.03
$\hat{\beta}_{2}(.)$ 0.01 0.01
$\hat{\beta}_{3}(.)$ 1 1
$\hat{\beta}_{4}(.)$ 0 0.01
$\hat{\beta}_{5}(.)$ 0.99 0.99
$\hat{\beta}_{6}(.)$ 0.01 0.02
Table 7: Proportion of non-zero estimates for each functional coefficient from the BGLSS method among the 100 replications, according to the data dispersion degree.
Concluding the comparative study, Figure 12 presents the estimates from the BGLSS method for the third and fifth functional coefficients. We can also observe that the BGLSS method leads to a smaller dispersion of the estimated functional coefficient curves when compared to the grLASSO, grSCAD and grMCP methods (see Figure 10). At the same time, it yields a dispersion that matches the levels of dispersion in the generated data.
(a) $\hat{\beta}_{3}(.)$ (BGLSS, $\sigma=0.2$).
(b) $\hat{\beta}_{3}(.)$ (BGLSS, $\sigma=20$).
(c) $\hat{\beta}_{5}(.)$ (BGLSS, $\sigma=0.2$).
(d) $\hat{\beta}_{5}(.)$ (BGLSS, $\sigma=20$).

### v5

@@ 3.3.1 Comparative analysis with grLASSO, grSCAD, grMCP and BGLSS
Among the frequentist methods that can be applied in this context of FOSR with functional coefficients expanded via basis functions, we can highlight the group Least Absolute Shrinkage and Selection Operator (grLASSO) (Yuan and Lin, 2006), the group Smoothly Clipped Absolute Deviation grSCAD) (Wang et al., 2007) and the group Minimax Concave Penalty (grMCP) (Huang et al., 2012). These methods are generalizations for group structures and derive from LASSO (Least Absolute Shrinkage and Selection Operator) (Tibshirani, 1996), SCAD (Smoothly Clipped Absolute Deviation) penalized regression (Fan and Li, 2001) and MCP (Minimax Concave Penalty) penalized regression (Zhang, 2010), which are regularization and variable selection methods for regression analysis.
Thus, the regularization and selection methods in groups described above (grLASSO, grSCAD and grMCP) are considered in this section to carry out a comparative analysis with our proposed method. To complete the comparisons with other existing methods, we also consider the Bayesian Group LASSO with Spike and Slab prior (BGLSS) method (Xu and Ghosh, 2015). In order to obtain the results of the grLASSO, grSCAD and grMCP models, we used the grpreg function from the grpreg package (Breheny et al., 2021) of the R statistical software, while the BGLSS function of the R package MBSGS (Liquet and Sutton, 2022) was used to implement BGLSS.
We used synthetic data replications generated as described in Subsection 3.1 and fitted our model using $K=10$ B-spline bases for each functional coefficient. Finally, it is worth noting that the models are compared in two scenarios, one simulating data with low dispersion ($\sigma=0.2$) and another with higher dispersion ($\sigma=20$). For each dispersion scenario, 100 simulated datasets were used. Implementing the competing methods requires building appropriate response vectors and design matrices. For the frequentist methods, a grid of values for the regularization parameter $\lambda$ is also needed. In Section 2 of the Supplementary Material, we provide all the details for implementing the competing methods.
@@
Figure 4: Boxplots of the metric (7), according to the model used and the data dispersion degree. The $1^{\text{st}}$ Proposed Model” is considered to be the version with $\mu$ as a parameter, while the $2^{\text{nd}}$ Proposed Model” is the version with $\mu$ as a hyperparameter ($\mu=0.3$ for $\sigma=0.2$ and $\mu=0.5$ for $\sigma=20$). For the competing methods grLASSO, grMCP, and grSCAD, the results correspond to the best $\lambda$ from the tested grid.
Supplementary Figures 8 and 9 show boxplots of metric (7) and MSE, respectively, for each competing model, grouped by data dispersion degree and tested $\lambda$ values. Regardless of the metric ((7) or MSE), the fit quality begins to decline beyond a certain $\lambda$, indicating that stronger regularization leads to suboptimal results when $\lambda$ becomes too large. Figure 4 and Supplementary Figure 10 compare our proposed model with grLASSO, grMCP, grSCAD, and BGLSS based on metric (7) and MSE, respectively, considering the best $\lambda$ from the tested grid for grLASSO, grMCP, and grSCAD.
@@
In order to demonstrate the capacity of each frequentist competing model to select the correct covariates across different values of $\lambda$ and dispersion levels, Supplementary Tables 1 and 2 present the proportion of times that each functional coefficient estimate was different from zero among the total of 100 replications. We notice that none of the models performed variable selection at the lowest regularization level ($\lambda=0.001$) across all three competing methods (all proportions were equal to 1). For the other $\lambda$ values tested, the models began selecting variables. Such selection for data with low dispersion is predominantly correct from $\lambda=0.250$, while for data with large dispersion from $\lambda=2.005$. However, the non-zero proportions in Supplementary Table 2 associated with the fifth functional coefficient become lower for larger values of $\lambda$.
@@
It is noted (Supplementary Table 2 and Supplementary Figure 9 (b)) that the competing models are not able to balance very well the accuracy in the selection in the scenario with greater data dispersion while preserving the quality of the fit according to the metrics under study, meaning that they need to lose fit quality and increase the regularization to gain better accuracy in the selection. The last finding from evaluating Supplementary Tables 1 and 2 is that the grSCAD and grMCP models stand out from the grLASSO model in the scenario with low data dispersion ($\sigma=0.2$), better selecting the variables for the different $\lambda$ values tested.
@@
Tables 2 and 3 allow a brief comparison regarding the capacity to select the correct covariates among the best configurations tested for the competing models, as well as for our proposed model and for the BGLSS. The best model configuration (for the frequentist methods) was chosen according to the performance metric (7). For this, the average metric value over the one hundred replications was calculated for each model configuration, and we then selected the model with the configuration that returned the highest of these averages.
@@
Table 2: Proportion of non-zero estimates for each functional coefficient ($\hat{Z}_{l}\hat{\beta}_{l}(.)$ for our proposed model and $\hat{\beta}_{l}(.)$ for the competing methods) among the 100 replications, according to the best tested model configuration. Data with $\sigma=0.2$.
@@
$\hat{Z}_{l}\hat{\beta}_{l}(.)$ or $\hat{\beta}_{l}(.)$ Proposed model ($\mu=0.3$ and $\lambda=\sqrt{2}$) Proposed model ($\mu$ as a parameter and $\lambda=\sqrt{2}$) grLASSO ($\lambda=$ $0.001$) grMCP ($\lambda=$ $0.250$) grSCAD ($\lambda=$ $0.250$) BGLSS
$l=1$ 0 0 1 0 0 0
$l=2$ 0 0 1 0 0 0.01
$l=3$ 1 1 1 1 1 1
$l=4$ 0 0 1 0 0 0
$l=5$ 1 1 1 1 1 0.99
$l=6$ 0 0 1 0 0 0.01
@@
Table 3: Proportion of non-zero estimates for each functional coefficient ($\hat{Z}_{l}\hat{\beta}_{l}(.)$ for our proposed model and $\hat{\beta}_{l}(.)$ for the competing methods) among the 100 replications, according to the best tested model configuration. Data with $\sigma=20$.
$\hat{Z}_{l}\hat{\beta}_{l}(.)$ or $\hat{\beta}_{l}(.)$ Proposed model ($\mu=0.5$ and $\lambda=\sqrt{2}$) Proposed model ($\mu$ as a parameter and $\lambda=\sqrt{2}$) grLASSO ($\lambda=$ $0.001$) grMCP ($\lambda=$ $0.500$) grSCAD ($\lambda=$ $0.001$) BGLSS
$l=1$ 0.04 0.01 1 0.95 1 0.03
$l=2$ 0.02 0.01 1 0.96 1 0.01
$l=3$ 1 1 1 1 1 1
$l=4$ 0.02 0 1 0.94 1 0.01
$l=5$ 1 0.99 1 1 1 0.99
$l=6$ 0.03 0.03 1 1 1 0.02
@@
The results presented in Table 2 show that most of the compared models correctly performed the covariate selection on low dispersion data ($\sigma=0.2$), except for grLASSO, which was unable to exclude any covariate in any of the hundred replications. When evaluating the selection performance for data with greater dispersion ($\sigma=20$), we can observe in Table 3 that only the two versions of the proposed model ($\mu$ as a hyperparameter and as a parameter) and the BGLSS were able to identify the two true variables that were used in the construction of synthetic data ($3^{\text{rd}}$ and $5^{\text{th}}$) in almost all replications.
@@
It is important to note that the value of $\lambda=\sqrt{2}$ used in the various versions of the proposed model represents the lowest feasible setting for this parameter. Using a smaller value can lead to problems with the uniqueness of the stationary distribution due to two factors: low regularization and multicollinearity among the covariates. While this setting offers the minimum possible regularization, the proposed model versions still achieved a good fit to the data while accurately selecting the variables. In contrast, as mentioned earlier, the grLASSO, grSCAD, and grMCP models faced difficulties in balancing accurate variable selection with the quality of fit.
@@
Finally, unlike the grLASSO, grSCAD and grMCP methods, the BGLSS method can select the variables well even in the case with large data dispersion ($\sigma=20$) and is the only method among the ones considered here that can compete with our proposed methodology in terms of predictive quality (through the metric (7) and MSE) and variable selection capacity.
Figure 4 and Supplementary Figure 10 present the results regarding the performance metric in (7) and MSE, respectively, for the best model configurations. The metric (7) was used as a reference for choosing the best model configurations to be compared in Figure 4, as it was done for Tables 2 and 3. Whereas, for Supplementary Figure 10, the MSE itself was the reference metric for defining the best model settings. Thus, the comparison via MSE becomes fair. Turning the analysis to the results from data with low dispersion ($\sigma=0.2$), regardless of the metric being analyzed, slightly better performance of the competing methods can be observed in comparison to the two versions of the proposed model, especially when using the metric (7) as a reference. As for the data with greater dispersion, the models present similar goodness-of-fit and estimation accuracy.
Note that for the BGLSS, in one replication with low dispersion ($\sigma=0.2$), the method had its worst performance, resulting in an outlier (for both evaluated metrics). Figure 4 and Supplementary Figure 10 omit this outlier since we adjusted the scale to focus on the interquartile range region to enable the comparison across the boxplots.
Finally, Figures 5 and 6 show the estimates obtained in each replication for the third and fifth functional coefficients for each method and each level of data dispersion. In this case, the performance metric in (7) was used as a reference to choose the model configurations to compare. When $\sigma=20$, it is noticeable that there is a smaller dispersion of the estimated functional coefficients obtained by our proposed model (first row of Figure 6) compared to the grLASSO, grMCP and grSCAD competing models. This observation is valid for both the third and the fifth functional coefficients.
(a) $\hat{\beta}_{3}(.)$ (proposed model, $\mu=0.3$).
(b) $\hat{\beta}_{5}(.)$ (proposed model, $\mu=0.3$).
(c) $\hat{\beta}_{3}(.)$ (proposed model, $\mu$ as a parameter).
(d) $\hat{\beta}_{5}(.)$ (proposed model, $\mu$ as a parameter).
(e) $\hat{\beta}_{3}(.)$ (grLASSO, $\lambda=0.001$).
(f) $\hat{\beta}_{5}(.)$ (grLASSO, $\lambda=0.001$).
(g) $\hat{\beta}_{3}(.)$ (grMCP, $\lambda=0.250$).
(h) $\hat{\beta}_{5}(.)$ (grMCP, $\lambda=0.250$).
(i) $\hat{\beta}_{3}(.)$ (grSCAD, $\lambda=0.250$).
(j) $\hat{\beta}_{5}(.)$ (grSCAD, $\lambda=0.250$).
(k) $\hat{\beta}_{3}(.)$ (BGLSS).
(l) $\hat{\beta}_{5}(.)$ (BGLSS).